What is the change of reaction when temperature decreases from 100°C to 70°c if the temperature coefficient is 2
1
Expert's answer
2020-10-08T13:59:08-0400

According to Van 't Hoff's empirical rule: 

vt+10 / vt = n(100-70)/10

vt+10 / vt = 23 = 8 times